A bridge is built in the shape of a semielliptical arch. It has a span of 94 feet. The height of the arch 28 feet from the center is to be 6 feet. Find the height of the arch at its center.
Answer by stanbon(48569)

A bridge is built in the shape of a semielliptical arch. It has a span of 94 feet. The height of the arch 28 feet from the center is to be 6 feet. Find the height of the arch at its center.
---------
Draw the picture putting the width of 94 on the x-axis centered
at the origin.
The vertices are then (-47,0) and (47,0)------
Then x^2/47^2 + y^2/b^2 = 1
The ellipse passes thru the point (28,6)
Substitute those values in the equation to solve for "b".
--------
28^2/47^2 + 6^2/b^2 = 1
0.3549117248 + 36/b^2 = 1
36/b^2 = 0.6450882752..
b^2/36=1.550175439
b^2=55.806
------
At the center the height is b=7.47 feet
===========
EQUATION of the ellipse
x^2/47^2 + y^2/55.806 = 1
==========
Cheers,
Stan H.